OpenStudy (emogirlatemoooocow):

n which measurement situations would the most appropriate unit of measurement be inches? Choose exactly two answers that are correct. A. width of the Mississippi River B. diameter of a pizza C. height of a skyscraper D. length of a pencil

OpenStudy (emogirlatemoooocow):

A rectangle has a length of 15 cm and a perimeter of 72 cm. Which equation could be used to find the width of the rectangle? A. 15w = 72 B. w + 15 = 72 C. 2w + 15 = 72 D. 2w + 30 = 72 this 1